OpenStudy (anonymous):

If sin α + cosec α = 2, then what does sin^n α + cosec^n α equal?

OpenStudy (anonymous):

The options provided are :- a) n(sinα+cosecα) b) 2 c) \[\frac{n}{\sin \alpha + cosec \alpha}\] d)None of these If the answer is being given from one of the options, kindly explain how the answer was obtained.

hartnn (hartnn):

do you know the relation between sin a and cosec a ?

hartnn (hartnn):

bdw, i am getting the answer as 2